OpenStudy (binary3i):

\[\sum_{}^{} [(f(x+dx,y) - f(x,y))\times(f(x,y+dy)-f(x,y))]\] will this give me the area of the graph f(x,y)?

OpenStudy (unklerhaukus):

shouldn't the \(\text dx\)'s be \(\Delta x\)'s?

OpenStudy (experimentx):

I bet it's a multivariate function f(x,y)

OpenStudy (anonymous):

i think it s a yes

OpenStudy (experimentx):

it seems like summation of \[ \sum f_x \Delta x\times f_y \Delta y \] should be \[ \sum \sum \Delta x \Delta y\]

OpenStudy (anonymous):

c graph is a function of x nd y....so take include z axis to.....now....for small variation of x nd y...... above equation is the sum of small squares......